A rancher has some sheep and ostriches.
His young daughter observed that the animals have a total
of 60 eyes and 86 feet. How many animals of each type
does the rancher have?

Answer by checkley77(12844)

2S+2O=60 total eyes. (MULTIPLY BY -2 & ADD.)
4S+2O=86 total feet.
-4S-4O=-120
----------------
-2O=-34
O=-34/-2
O=17 OSTRICHES.
2S+2*17=60
2S+34=60
2S=60-34
2S=26
S=26/2
S=13 SHEEP.
PROOF:
4*13+2*17=86
52+34=86
86=86